Help I need help answering this question, Im in college Algebra!
Suppose that you have a supply of a 35% solution of alcohol and a 65% solution of alcohol. How many quarts of each should be mixed to produce 30 quarts that is 60% alcohol?
35% solution=___________qt
65% solution=___________qt

Answer by stanbon(75887)

Suppose that you have a supply of a 35% solution of alcohol and a 65% solution of alcohol. How many quarts of each should be mixed to produce 30 quarts that is 60% alcohol?
-----------------------
Alcohol + alcohol = alcohol
0.35x + 0.65(30-x) = 0.50*30
---
Multiply thru by 100 and solve for "x":
35x + 65*30 - 65x = 50*30
-35x = -15*30
x = (3/7)30
x = 12.86 quarts ; (amt. of 35% solution needed)
30 - 12.86 = 17.14 quarts ; (amt of 65% solution needed)
==========================
Cheers,
Stan H.
==========================